OpenStudy (anonymous):

Find the length of the diameter of a sphere with a volume of 143.8 miles^3.

OpenStudy (anonymous):

5.86 miles 3.25 miles 6.5 miles 11.72 miles

OpenStudy (ajprincess):

The volume of the sphere=(4/3)*pi*r^3. r-radius. d-diameter d=2r (4/3)*pi*r^3=143.8 find r from. then plug in that value of r in d=2r to find d. Does that help?

OpenStudy (anonymous):

ok how do I plug in for R

OpenStudy (ajprincess):

(4/3)*pi*r^3=143.8 from this can u find r?

OpenStudy (anonymous):

No need help

OpenStudy (ajprincess):

\[\frac{ 4 }{ 3 }*\pi*r^3=143.8\] \[r^3=\frac{ 143.8*3 }{ 4*\pi }\] \[=?\]

OpenStudy (anonymous):

6.5 Miles

OpenStudy (anonymous):

??

OpenStudy (ajprincess):

yup:)